The denominator of the fraction is 1 less than twice the numerator.If the numerator & denominator are increased both by 1, the fraction becomes 3/5.Find the fraction.
(A)5/9
(B)3/5
(C)4/7

Answer by richwmiller(17219)

a) 2*5-1=9
ok
6/10=3/5
ok
answer a) 5/9